(5x^2 -6)^(1/4) =x
Thank you,
Alexus

Answer by stanbon(75887)

Raise both sides to the 4th power, as follows:
5x^2-6=x^4
Rearrange:
x^4-5x^2+6=0
Factor:
(x^2-3)(X^2-2)=0
X^2=3 OR X^2=2
X=sqrt3 or x=-sqrt3 or x=sqrt2 or x=-sqrt2
Cheers,
stan H.